Which is larger, the volume of the prism or the volume of the pyramid?
The prism cube is 10 m x 10 m x 10 m.
The pyramid is a square pyramid.
The square Is 10 m x 10 m and it is 10 m high.

Answer by jim_thompson5910(35256)

The prism (which is more specifically a cube in this case) is larger by volume

Volume of cube = 10*10*10 = 1,000 cubic meters

Volume of square pyramid = (1/3)*10*10*10 = 1000/3 = 333.33 cubic meters (approximately)

Essentially, we can fit 3 of these pyramids into 1 cube assuming we can cut the pyramid up into smaller pieces. The formulas I used are:

Volume of rectangular prism = (length)*(width)*(height)
volume of rectangular pyramid = (1/3)*(length)*(width)*(height)